# ZeroScene: A Zero-Shot Framework for 3D Scene Generation from a Single Image and Controllable Texture Editing

### v1

@@ Abstract
In the field of 3D content generation, single image scene reconstruction methods still struggle to simultaneously ensure the quality of individual assets and the coherence of the overall scene in complex environments, while texture editing techniques often fail to maintain both local continuity and multi-view consistency. In this paper, we propose a novel system ZeroScene, which leverages the prior knowledge of large vision models to accomplish both single image-to-3D scene reconstruction and texture editing in a zero-shot manner. ZeroScene extracts object-level 2D segmentation and depth information from input images to infer spatial relationships within the scene. It then jointly optimizes 3D and 2D projection losses of the point cloud to update object poses for precise scene alignment, ultimately constructing a coherent and complete 3D scene that encompasses both foreground and background. Moreover, ZeroScene supports texture editing of objects in the scene. By imposing constraints on the diffusion model and introducing a mask-guided progressive image generation strategy, we effectively maintain texture consistency across multiple viewpoints and further enhance the realism of rendered results through Physically Based Rendering (PBR) material estimation. Experimental results demonstrate that our framework not only ensures the geometric and appearance accuracy of generated assets, but also faithfully reconstructs scene layouts and produces highly detailed textures that closely align with text prompts. Leveraging generative artificial intelligence, ZeroScene can transform 2D images into diversified 3D worlds with various styles, showing broad application potential in virtual content creation, such as digital twins and immersive game production, while also effectively support ”real-to-sim” transfer in robotics through the generation of highly realistic and trainable simulation environments. Project page: https://xdlbw.github.io/ZeroScene/
Figure 1: ZeroScene is capable of generating independent foreground instances while employing dedicated background modeling, thereby reconstructing a complete scene that is spatially precisely aligned with the input image. Moreover, it supports high-fidelity texture editing for objects within the scene and enables physically based rendering.
22footnotetext: 🖂 Corresponding author.
@@ 1 Introduction
3D generation technology enables the creation of immersive and realistic objects and environments, significantly enhancing the intuitiveness and interactive experience of digital content. As an important research direction in computer graphics, it holds substantial application value in fields such as digital twins, virtual reality, and embodied intelligence. When we focus on the specific task of single RGB image-to-3D generation, although current methods [55, 45, 47, 28] have demonstrated remarkable performance in generating single objects, their quality still declines significantly when handling complex scenes containing multiple objects. This is primarily due to mutual occlusion and overlap between objects in the single view, leading to loss of detail and multi-view inconsistency in the generated assets. Furthermore, the lack of modeling of spatial relationships between objects often results in physically implausible generated scenes, manifested as abnormal object poses, mutual penetration, floating, or failure to form reasonable support relationships under physical contact constraints. Although several recent works have explored compositional scene synthesis, they still commonly suffer from limitations such as insufficient understanding of spatial relationships [17, 57] and poor geometric and texture quality of generated objects [9, 1]. Moreover, only a limited number of studies [1, 5] have involved the background components within scenes, and exploration in this aspect remains relatively preliminary.
In addition to the paradigm of jointly generating object shapes and textures, another research route focuses on texture editing based on given 3D geometries. Detailed and photorealistic textures are crucial for enhancing the visual appearance of 3D objects, and efficient, high-quality texture synthesis and editing techniques will further advance industries such as animation and gaming. In recent years, with the rapid development of 2D image generation techniques [38, 2], text or image-guided texture generation methods [3, 48] have made significant progress. Many approaches [36, 7, 44] leverage pre-trained image diffusion models [37, 40] by incorporating depth constraints to guide the synthesis of RGB images that conform to geometric surfaces. However, these methods often suffer from limited generation quality and multi-view inconsistency, resulting in blurriness or artifacts in the finally synthesized texture. SyncMVD [30] synchronizes multi-view information at each denoising step and employs cross-view attention to facilitate information sharing, thereby generating detailed and seamless textures. Nevertheless, this approach still struggles to completely avoid the ”multi-face janus problem”. Therefore, achieving high-quality texture generation that is globally consistent across views, seamlessly aligned, and faithful to user intent remains a significant challenge in current research.
@@ 2.1 Single Image Scene Generation
Generating 3D scenes from a single image is a challenging task, with difficulties including handling object diversity, addressing severe occlusions, and accurately understanding spatial relationships among various parts of the scene. Fortunately, a variety of mature vision models are widely used in related subtasks, such as image segmentation [26, 29, 35], image inpainting [39, 25], 3D reconstruction [42, 49], and 3D content generation [51, 46, 27]. Existing methods often rely on these foundational models, integrating decomposition-recomposition [17, 9, 57, 1, 8], scene graph construction [12], optimization strategies [17, 9, 50], and point cloud alignment [57, 1] to achieve multi-asset generation and scene layout recovery.
Recent studies have attempted to more directly model spatial relationships between objects. For instance, CAST [50] proposes a component-aligned generation strategy that aligns content generated in canonical space with scene space by computing model transformation matrices, enabling efficient and accurate 3D scene generation from a single image. MIDI [22] captures spatial dependencies during the diffusion process through a multi-instance attention mechanism, yet it still faces certain limitations when handling complex object interactions. Similarly, Scenegen [31] employs local-global attention blocks to facilitate interaction among different assets, but its generated results often exhibit redundant structures and artifacts.
@@ 2.2 Texture Editing
Recently, the research focus of AI-based 3D texture generation and editing has gradually shifted from generative adversarial networks [16] to diffusion models [20, 37]. This transition has enhanced the visual diversity and realism of textures, overcoming the limitations in flexibility and efficiency inherent in traditional manual creation methods. Methods based on iterative rendering [36, 33, 44] leverage depth or normal maps and employ ControlNet [54] to guide diffusion models, effectively aligning textures with geometric models. However, their outputs often suffer from insufficient detail clarity and visual consistency. Mapa [56] decomposes a 3D mesh into multiple segments and employs segment-controlled strategy to produce 2D images aligned with different mesh components. Several approaches build upon multi-view images [7, 24, 53, 23], incorporating UV inpainting modules to complete missing texture regions and achieve more comprehensive texture synthesis. MVPaint [10] addresses common issues in existing text-to-texture methods, such as local discontinuities, view inconsistency, and over reliance on UV unwrapping quality, through modules for multi-view consistent generation, inpainting of unobserved regions, and UV optimization. Nevertheless, its underlying multi-view diffusion model struggles to comprehend fine-grained information in user inputs, resulting in generated textures that lack detail. The latest research aims to produce textures with physical realism, particularly in terms of light response and material attributes. Paint-it [52] and MaterialMVP [18] are capable of generating a full set of texture maps, including albedo, roughness, metallic, and normal maps, supporting high resolution PBR pipelines. Material Anything [21] introduces a confidence mask mechanism to dynamically adjust the influence of illumination, thereby ensuring physical plausibility of material appearance under varying lighting conditions.
@@ 3.1 Foreground Object Generation and Composition
Instance Segmentation and Generation. Given a single image, our framework first performs foreground-background decoupling, followed by foreground object detection [29]. Using predefined semantic labels, it identifies bounding boxes, category labels, and confidence scores for each instance. Subsequently, based on the detected bounding boxes, a refined instance segmentation module [35] is employed to perform pixel-level extraction on $N$ candidate foreground regions, producing corresponding segmented images and binary masks $\{{p_{i},m_{i}}\}_{i=1}^{N}$. Due to mutual occlusion between instances in the original image, structural incompleteness often occurs. Directly using such incomplete instances for 3D generation would lead to degraded model quality. To address this, we leverage the powerful semantic reasoning capability of Vision-Language Models (VLMs), e.g., GPT-4o [25] to locate missing regions via text prompts and generate structurally complete inpainted images. To align with the rapidly advancing 3D generation techniques in the community, our system employs open-source models for image-to-3D conversion, thereby ensuring extensibility and flexible compatibility with future model advancements. Specifically, we use Hunyuan3D 2.5 [27] to map the inpainted single view instance images into 3D models $\{\mathcal{M}_{i}\}_{i=1}^{N}$. The generated 3D assets exhibit accurate geometry, high-fidelity texture consistency with the input image, and are equipped with PBR materials, resulting in high visual realism and rendering quality.
Point Cloud Extraction. In the second stage, we focus on recovering a 3D point cloud representation of the scene from the images, thereby establishing a geometric foundation for subsequent object layout optimization. DUSt3R [42] significantly simplifies this process. Trained on a large-scale dataset comprising 8.5 million image pairs, including indoor, outdoor, object-level, and scene-level samples, this model efficiently estimates depth information and camera parameters without relying on camera calibration. It constructs dense mappings between 2D pixels and 3D points, ultimately outputting the 3D structure in the form of pointmaps. Based on the obtained pointmaps, we further extract the complete scene point cloud $\mathcal{PC}$ and its corresponding camera parameters $\mathcal{C}$. Subsequently, we combine the scene point cloud $\mathcal{PC}$ with the instance-level masks $\{{m_{i}}\}_{i=1}^{N}$ obtained from the previous stage to perform spatial segmentation, thereby generating independent point cloud representations $\{\mathcal{PC}_{i}\}_{i=1}^{N}$ for each of the detected $N$ instances.
@@
Layout Optimization. Finally, ZeroScene strictly adheres to the scene layout of the original image to refine the position and pose of objects in 3D space. Concretely, we reconstruct point clouds $\mathcal{M}$ from the surface of each generated object and parameterize its pose as a set of learnable spatial transformation parameters $\phi=\{T,R,S\}$, denoting translation, rotation, and isotropic scaling, respectively. These parameters are optimized via gradient descent to achieve geometric alignment between the 3D objects and the image scene. The optimization objective is to minimize the 3D Chamfer Distance (CD) loss, which enforces consistency between the instance point cloud $\mathcal{M}_{i}$ and the target point cloud $\mathcal{PC}_{i}$ in 3D space. However, relying solely on 3D constraints can lead to instability during the translation and rotation optimization. Inspired by previous studies [6, 57], we introduce an additional 2D projection constraint as auxiliary supervision: using known camera parameters $\mathcal{C}$, we project both the instance point cloud $\mathcal{M}_{i}$ and the target point cloud $\mathcal{PC}_{i}$ onto the image plane, constructing a 2D CD loss that forces the two projected point sets to match, thereby facilitating 3D structure learning and enhancing cross-dimensional geometric consistency. The overall loss function is defined as:
$\begin{split}\mathcal{L}&=\lambda_{1}\cdot\mathcal{L}_{CD}^{3D}(\mathcal{M}_{i},\mathcal{PC}_{i})\\ &+\lambda_{2}\cdot\mathcal{L}_{CD}^{2D}(\text{Proj}_{\mathcal{C}}(\mathcal{M}_{i}),\text{Proj}_{\mathcal{C}}(\mathcal{PC}_{i}))\end{split}$ (1)
where $\lambda_{1}$ and $\lambda_{2}$ are weighting coefficients balancing the different supervision signals, and $\text{Proj}_{\mathcal{C}}(\cdot)$ denotes the projection operation based on the camera projection matrix. Fig. 3 illustrates the optimization process of point cloud registration in both 3D and 2D spaces.
@@ 3.2 Background Handling
Considering that large portions of the background are frequently occluded by foreground objects, the corresponding point cloud data is incomplete. To address this, we employ a VLM [25] for iterative object removal, obtaining a complete background image devoid of all foreground instances and their shadows. Based on this image, we utilize DUSt3R [42] to re-estimate a colored background point cloud. Subsequently, through post-processing steps such as denoising, vertex normal estimation, Poisson reconstruction, and planar smoothing, we obtain a colored explicit geometric model of the background. Finally, we employ Hunyuan3D-Paint 2.1 [14, 18] to generate PBR material textures. Similar to foreground instances, we represent background entities as a set of learnable parameters $\phi=\{T,R,S\}$ in 3D space and optimize their layout using the loss function in Eq. 1, thereby achieving a unified and harmonious assembly of the complete scene with both foreground and background elements.

### v2

@@ Abstract
In the field of 3D content generation, single image scene reconstruction methods still struggle to simultaneously ensure the quality of individual assets and the coherence of the overall scene in complex environments, while texture editing techniques often fail to maintain both local continuity and multi-view consistency. In this paper, we propose a novel system ZeroScene, which leverages the prior knowledge of large vision models to accomplish both single image-to-3D scene reconstruction and texture editing in a zero-shot manner. ZeroScene extracts object-level 2D segmentation and depth information from input images to infer spatial relationships within the scene. It then jointly optimizes 3D and 2D projection losses of the point cloud to update object poses for precise scene alignment, ultimately constructing a coherent and complete 3D scene that encompasses both foreground and background. Moreover, ZeroScene supports texture editing of objects in the scene. By imposing constraints on the diffusion model and introducing a mask-guided progressive image generation strategy, we effectively maintain texture consistency across multiple viewpoints and further enhance the realism of rendered results through Physically Based Rendering (PBR) material estimation. Experimental results demonstrate that our framework not only ensures the geometric and appearance accuracy of generated assets, but also faithfully reconstructs scene layouts and produces highly detailed textures that closely align with text prompts. Leveraging generative artificial intelligence, ZeroScene can transform 2D images into diversified 3D worlds with various styles, showing broad application potential in virtual content creation, such as digital twins and immersive game production, while also effectively supports "real-to-sim" transfer in robotics through the generation of highly realistic and trainable simulation environments. Project page: https://xdlbw.github.io/ZeroScene.
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371</concept_id> <concept_desc>Computing methodologies Computer graphics</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010178</concept_id> <concept_desc>Computing methodologies Artificial intelligence</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
[…]
volume: 45issue: 2
@@ 1 Introduction
3D generation technology enables the creation of immersive and realistic objects and environments, significantly enhancing the intuitiveness and interactive experience of digital content. As an important research direction in computer graphics, it holds substantial application value in fields such as digital twins, virtual reality, and embodied intelligence. When we focus on the specific task of single RGB image-to-3D generation, although current methods [ZWZ24, WLC24b, XLX25, LZL25b] have demonstrated remarkable performance in generating single objects, their quality still declines significantly when handling complex scenes containing multiple objects. This is primarily due to mutual occlusion and overlap between objects in the single view, leading to loss of detail and multi-view inconsistency in the generated assets. Furthermore, the lack of modeling of spatial relationships between objects often results in physically implausible generated scenes, manifested as abnormal object poses, mutual penetration, floating, or failure to form reasonable support relationships under physical contact constraints. Although several recent works have explored compositional scene synthesis, they still commonly suffer from limitations such as insufficient understanding of spatial relationships [HYL25, ZLH24] and poor geometric and texture quality of generated objects [CWW24, AÖE25]. Moreover, only a limited number of studies [AÖE25, CJL25] have involved the background components within scenes, and exploration in this aspect remains relatively preliminary.
In addition to the paradigm of jointly generating object shapes and textures, another research route focuses on texture editing based on given 3D geometries. Detailed and photorealistic textures are crucial for enhancing the visual appearance of 3D objects, and efficient, high-quality texture synthesis and editing techniques will further advance industries such as animation and gaming. In recent years, with the rapid development of 2D image generation techniques [SCS22, BBB25], text or image-guided texture generation methods [BKA24, XGF25] have made significant progress. Many approaches [RMA23, CSL23, WLC25] leverage pre-trained image diffusion models [RBL22, Tea24] by incorporating depth constraints to guide the synthesis of RGB images that conform to geometric surfaces. However, these methods often suffer from limited generation quality and multi-view inconsistency, resulting in blurriness or artifacts in the finally synthesized texture. SyncMVD [LXLW24] synchronizes multi-view information at each denoising step and employs cross-view attention to facilitate information sharing, thereby generating detailed and seamless textures. Nevertheless, this approach still struggles to completely avoid the "multi-face janus problem". Therefore, achieving high-quality texture generation that is globally consistent across views, seamlessly aligned, and faithful to user intent remains a significant challenge in current research.
@@ 2.1 Single Image Scene Generation
Generating 3D scenes from a single image is a challenging task, with difficulties including handling object diversity, addressing severe occlusions, and accurately understanding spatial relationships among various parts of the scene. Fortunately, a variety of mature vision models are widely used in related subtasks, such as image segmentation [KMR23, LZR24, RLZ24], image inpainting [SLM22, HLG24], 3D reconstruction [WLC24a, YSL25], and 3D content generation [YWL25, WLZ25, LZL25a]. Existing methods often rely on these foundational models, integrating decomposition-recomposition [CWW24, ZLH24, DWJ24, HYL25, WIR25, AÖE25, CWS25], Large Language Models [LZZZ24, GCL25], scene graph construction [DYY25], optimization strategies [CWW24, HYL25, GCL25, WIR25, YZY25], and point cloud alignment [ZLH24, AÖE25] to achieve multi-asset generation and scene layout recovery. Recent studies have attempted to more directly model spatial relationships between objects. For instance, CAST [YZY25] proposes a component-aligned generation strategy that aligns content generated in canonical space with scene space by computing model transformation matrices, enabling efficient and accurate 3D scene generation from a single image. MIDI [HGA25] captures spatial dependencies during the diffusion process through a multi-instance attention mechanism and Scenegen [MWZX26] employs local-global attention blocks to facilitate interaction among different assets.
Distinct from the component alignment adopted by CAST or the strategy proposed by Zhou et al. [ZLH24], which calculates the loss by randomly sampling discrete points within the contours of original instance masks, we achieve accurate layout recovery by minimizing the joint 3D and 2D projection loss between the predicted instance point clouds and the target point clouds. This effectively rectifies the scale/depth ambiguity inherent in single-view depth estimation, ensuring that the generated models strictly adhere to the spatial layout of the input images. Furthermore, in contrast with methods such as MIDI and Scenegen, which often suffer from object misalignment, structural incompleteness, or redundant artifacts due to the lack of explicit handling of occluded regions, our workflow fundamentally guarantees the structural integrity of occluded objects and the quality of individual assets by separating foreground from background and performing targeted image inpainting prior to 3D lifting.
@@ 2.2 Texture Editing
Recently, the research focus of AI-based 3D texture generation and editing has gradually shifted from generative adversarial networks [GPAM14] to diffusion models [HJA20, RBL22]. This transition has enhanced the visual diversity and realism of textures, overcoming the limitations in flexibility and efficiency inherent in traditional manual creation methods. Methods based on iterative rendering [RMA23, PWMAZ24, WLC25] leverage depth or normal maps and employ ControlNet [ZRA23] to guide diffusion models, effectively aligning textures with geometric models. However, their outputs often suffer from insufficient detail clarity and visual consistency. Mapa [ZPX24] decomposes a 3D mesh into multiple segments and employs segment-controlled strategy to produce 2D images aligned with different mesh components. Several approaches build upon multi-view images [CSL23, HGZ24, ZCQ24, HGW25], incorporating UV inpainting modules to complete missing texture regions and achieve more comprehensive texture synthesis. MVPaint [CMZ25] addresses common issues in existing text-to-texture methods, such as local discontinuities, view inconsistency, and over reliance on UV unwrapping quality, through modules for multi-view consistent generation, inpainting of unobserved regions, and UV optimization. Nevertheless, its underlying multi-view diffusion model struggles to comprehend fine-grained information in user inputs, resulting in generated textures that lack detail. The latest research aims to produce textures with physical realism, particularly in terms of light response and material attributes. Paint-it [YOPM24] and MaterialMVP [HYY25] are capable of generating a full set of texture maps, including albedo, roughness, metallic, and normal maps, supporting high resolution PBR pipelines. Material Anything [HWLW25] introduces a confidence mask mechanism to dynamically adjust the influence of illumination, thereby ensuring physical plausibility of material appearance under varying lighting conditions.
Methods like TEXTure [RMA23] rely excessively on depth-guided image inpainting, which tends to cause seam inconsistencies or visual artifacts when processing the surfaces of complex geometric objects. To address such limitations, we significantly enhance the geometric alignment and visual continuity of textures on complex topological structures, via incorporating richer geometric priors and combining the mask-guided progressive multi-view image generation strategy with a confidence-based back projection module. Moreover, our work not only focuses on the color generation of textures, but also integrates PBR material estimation, thereby endowing the generated textured meshes with favorable physical realism.
@@ 3.1 Foreground Object Generation and Composition
Instance Segmentation and Generation. Given a single image, our framework first performs foreground-background decoupling, followed by foreground object detection [LZR24]. Using predefined semantic labels, it identifies bounding boxes, category labels, and confidence scores for each instance. Subsequently, based on the detected bounding boxes, a refined instance segmentation module [RLZ24] is employed to perform pixel-level extraction on $N$ candidate foreground regions, producing corresponding segmented images and binary masks $\{{p_{i},m_{i}}\}_{i=1}^{N}$. Due to mutual occlusion between instances in the original image, structural incompleteness often occurs. Directly using such incomplete instances for 3D generation would lead to degraded model quality. To address this, we leverage the powerful semantic reasoning capability of Vision-Language Models (VLMs), e.g., GPT-4o [HLG24] to locate missing regions via text prompts and generate structurally complete inpainted images. To align with the rapidly advancing 3D generation techniques in the community, our system employs open-source models for image-to-3D conversion, thereby ensuring extensibility and flexible compatibility with future model advancements. Specifically, we use Hunyuan3D 2.5 [LZL25a] to map the inpainted single view instance images into 3D models $\{\mathcal{M}_{i}\}_{i=1}^{N}$. The generated 3D assets exhibit accurate geometry, high-fidelity texture consistency with the input image, and are equipped with PBR materials, resulting in high visual realism and rendering quality.
@@
Point Cloud Extraction. In the second stage, we focus on recovering a 3D point cloud representation of the scene from the images, thereby establishing a geometric foundation for subsequent object layout optimization. DUSt3R [WLC24a] significantly simplifies this process. Trained on a large-scale dataset comprising 8.5 million image pairs, including indoor, outdoor, object-level, and scene-level samples, this model efficiently estimates depth information and camera parameters without relying on camera calibration. It constructs dense mappings between 2D pixels and 3D points, ultimately outputting the 3D structure in the form of pointmaps. Based on the obtained pointmaps, we further extract the complete scene point cloud $\mathcal{PC}$ in the camera coordinate system and its corresponding camera parameters $\mathcal{C}$. Subsequently, we combine the scene point cloud $\mathcal{PC}$ with the instance-level masks $\{{m_{i}}\}_{i=1}^{N}$ obtained from the previous stage to perform spatial segmentation, thereby generating independent point cloud representations $\{\mathcal{PC}_{i}\}_{i=1}^{N}$ for each of the detected $N$ instances.
Layout Optimization. Finally, ZeroScene strictly adheres to the scene layout of the original image to refine the position and pose of objects in 3D space. Concretely, we reconstruct point clouds $\mathcal{M}$ from the surface of each generated object and parameterize its pose as a set of learnable spatial transformation parameters $\phi=\{T,R,S\}$, denoting translation, rotation, and isotropic scaling, respectively. These parameters are optimized via gradient descent to achieve geometric alignment between the 3D objects and the image scene. The optimization objective is to minimize the 3D Chamfer Distance (CD) loss, which enforces consistency between the instance point cloud $\mathcal{M}_{i}$ and the target point cloud $\mathcal{PC}_{i}$ in 3D space. However, relying solely on 3D constraints can lead to instability during the translation and rotation optimization. Inspired by previous studies [CHLZ21, ZLH24], we introduce an additional 2D projection constraint as auxiliary supervision: using known camera parameters $\mathcal{C}$, we project both the instance point cloud $\mathcal{M}_{i}$ and the target point cloud $\mathcal{PC}_{i}$ onto the image plane, constructing a 2D CD loss that forces the two projected point sets to match, thereby facilitating 3D structure learning and enhancing cross-dimensional geometric consistency. The overall loss function is defined as:
$\mathcal{L}=\lambda_{1}\cdot\mathcal{L}_{CD}^{3D}(\mathcal{M}_{i},\mathcal{PC}_{i})+\lambda_{2}\cdot\mathcal{L}_{CD}^{2D}(\text{Proj}_{\mathcal{C}}(\mathcal{M}_{i}),\text{Proj}_{\mathcal{C}}(\mathcal{PC}_{i}))$ (1)
where $\lambda_{1}$ and $\lambda_{2}$ are weighting coefficients balancing the different supervision signals, and $\text{Proj}_{\mathcal{C}}(\cdot)$ denotes the perspective projection operation based on the camera projection matrix. Fig. 2 illustrates the optimization process of point cloud registration in both 3D and 2D spaces.
@@ 3.2 Background Handling
Considering that large portions of the background are frequently occluded by foreground objects, the corresponding point cloud data is incomplete. To address this, we employ a VLM [HLG24] for iterative object removal, obtaining a complete background image devoid of all foreground instances and their shadows. Based on this image, we utilize DUSt3R [WLC24a] to re-estimate a colored background point cloud. Since the background is generally composed of regular planar surfaces such as floors and walls, we first partition the point cloud into distinct regions according to its normal distribution. For the point cloud subset $\mathcal{B}=\{b_{i}\}_{i=1}^{N}$ corresponding to each region, we adopt robust RANSAC sampling combined with the constrained least squares method to extract its principal plane parameters $\mathbf{\pi}=[\mathbf{n}^{\top},d]^{\top}$, where $\mathbf{n}\in\mathbb{R}^{3}$ denotes the unit normal vector and $d$ is the intercept. Specifically, we optimize the parameters by minimizing the sum of squared distances from the inlier set $\mathcal{S}\subseteq\mathcal{B}$ to the plane, formulated as:
$\min_{\mathbf{n},d}\sum_{b_{i}\in\mathcal{S}}\|\mathbf{n}^{\top}b_{i}+d\|^{2},\quad\text{s.t.}\ \|\mathbf{n}\|_{2}=1$ (2)
Subsequently, we perform denoising on the point cloud with the constraints imposed by the fitted plane, and take this as a geometric prior to guide the subsequent Poisson reconstruction, thus yielding a smooth and continuous background mesh with color attributes. Finally, we perform scale calibration on the background model according to the known foreground depth range, and employ Hunyuan3D-Paint 2.1 [FYY25, HYY25] to generate PBR material textures. Similar to foreground instances, we represent background entities as a set of learnable parameters $\phi=\{T,R,S\}$ in 3D space and optimize their layout using the loss function in Eq. 1, thereby achieving a unified and harmonious assembly of the complete scene with both foreground and background elements.
Figure 4: Mask-guided progressive image generation strategy. "Project" refers to the process of projecting all currently known regions into the latent space of the next viewpoint. "DM" denotes a pre-trained text-to-image diffusion model [Tea24] equipped with ControlNet [ZRA23].